OpenStudy (countrygirl1431):

Solve the following equation. x + 5 = 2x

pooja195 (pooja195):

Subtracy x from both sides 2-1=?

OpenStudy (anonymous):

5

pooja195 (pooja195):

bruh

OpenStudy (countrygirl1431):

4 - 7x + 10 = 15 - 6x - 8

OpenStudy (anonymous):

x=2

pooja195 (pooja195):

Really?